OpenStudy (anonymous):

what is Log 2 ^3/Log 2 ^4 ? the same as?

OpenStudy (anonymous):

3/4?

OpenStudy (anonymous):

log2^3 = 3 * log2 log2^4 = 4 * log2 ===> (3*log2) / (4*log2) = 3/4.